# CodeReviewQA: The Code Review Comprehension Assessment for Large Language Models

### v1

@@ Abstract
State-of-the-art large language models (LLMs) have demonstrated impressive code generation capabilities but struggle with real-world software engineering tasks, such as revising source code to address code reviews, hindering their practical use. Code review comments are often implicit, ambiguous, and colloquial, requiring models to grasp both code and human intent. This challenge calls for evaluating large language models’ ability to bridge both technical and conversational contexts. While existing work has employed the automated code refinement (ACR) task to resolve these comments, current evaluation methods fall short, relying on text matching metrics that provide limited insight into model failures and remain susceptible to training data contamination. To address these limitations, we introduce a novel evaluation benchmark, CodeReviewQA that enables us to conduct fine-grained assessment of model capabilities and mitigate data contamination risks. In CodeReviewQA, we decompose the generation task of code refinement into three essential reasoning steps: change type recognition (CTR), change localisation (CL), and solution identification (SI). Each step is reformulated as multiple-choice questions with varied difficulty levels, enabling precise assessment of model capabilities, while mitigating data contamination risks. Our comprehensive evaluation spans 72 recently released large language models on 900 manually curated, high-quality examples across nine programming languages. Our results show that CodeReviewQA is able to expose specific model weaknesses in code review comprehension, disentangled from their generative automated code refinement results.  [^1] All code and data will be released upon acceptance.
@@ 5.2 Model Selection Criteria
In total, we selected 72 state-of-the-art open source LLMs, that have satisfied the three criteria. The included models are considered state-of-the-art as of March, 2025. See Table 6 in Appendix E for descriptions of all 72 models. The models are grouped into five scales based on their model parameters: $\leq$3B, $\leq$9B, $\leq$16B, $\leq$34B, and $\leq$72B. We selected models with $\leq$72B parameters, as it is the largest non-quantised model class that we can support locally to extract answer probabilities.
5.3 Implementation Details
The experiments were carried out on a single node consisting of 64 cores (Intel(R) Xeon(R) Platinum 8462Y+ @ 2.80GHz), 928GBs of RAM and four GPUs (NVIDIA H100-80GB SXM5). For efficient inference, we utilised the default vLLM implementation with Hugging Face models  [^3] Models were imported from Hugging Face Model Hub. For all tasks, we set the temperature to zero for greedy decoding.
@@ 6 Results Overview
To compare how models of different scales perform on CodeReviewQA, we conducted experiments using all 72 models. Due to space limitations, complete results on ACR and all three MCQA probes are presented in Appendix E. In Table 6, we compare the MCQA probe results of the top-2 models in terms of ACR from each scale class.
@@ 7 Insights from Probing ACR Failures
The results are presented in Table 7, including the percentage of failure on each of the MCQA probes, [^4] For CI and SI, a failure in either difficulties is counted. and the overall failure on at least one probe ($\geq 1$ Probe). We find that all five models failed at least one MCQA probe for 76.5%-99.8% of the failure cases, indicating that the model struggled at the code review comprehension stage. CTR failures are seldom associated with the ACR failures, as this capability only accounts for 22.2%-37.4% of cases. In contrast, CL failures are often associated with ACR failures, particularly for the three smaller models, where 74.0%-99.4% of the cases overlap with ACR failures. SI failures account for 95.8% of cases for Qwen2.5-Coder-3B-Instruct. Thus, for the smallest model, most non-exact matches are associated with failures in both CL and SI. For both Qwen2.5-Coder-7B-Instruct and Qwen2.5-Coder-14B-Instruct, CL failure is the major factor associated with ACR failure. For the larger models, gemma-2-27b-it and Llama-3.1-70B-Instruct, the probe failures are more evenly distributed, varying based on the model’s specific incompetencies.
@@
Table 3: MCQA Failure (%) in Non-Exact Match Cases of Top Performing Models (per scale class) in ACR.
@@ 8 Evaluating Data Contamination
$\text{PPL}(\textbf{X})=\exp(-\frac{1}{t}\sum^{t}_{t=0}\log p_{\theta}(x_{i}|x_{<i})),$ (5)
@@
$\text{NG}(\textbf{X})=\frac{1}{\eta\cdot K}\sum^{\eta}_{i=0}\sum^{K}_{j=0}I(X_{s_{j}:s_{j}+n},\hat{X}_{s_{j}:s_{j}+n}),$ (6)
@@ Appendix D Descriptive Statistics
Specification ratio is the code edit distance of the change divided by the length of its respective code review comment. It is calculated as Code edit distanceComment $\frac{\text{Code edit distance}}{\text{Comment length}}$, and can be interpreted as the number of character edits in the code change with respect to each word in the comment. Since code review comments may be under-specified and implicit, we use specification ratio as a heuristic metric that incorporates this notion. Intuitively, examples with larger code edit distances and shorter comment lengths i.e. larger specification ratios, may be under-specified in its description of the required code change. Table D shows that the average specification ratio is 4.88. See Figure 7 for examples of different specification ratios.
@@
Figure 2: Prompt Templates.
[…]
Figure 3: Examples of Variation in Difficulty.
Figure 4: Examples of Different Comment Lengths.
Figure 5: Examples of Different Code Edit Distances.
Figure 6: Examples of Different Code Element Ratios.
Figure 7: Examples of Different Specification Ratios.
Appendix E Experimental Results on 72 models
This section shows our complete results. The details of all evaluated LLMs are presented in Table 6, and their achieved exact match rate on ACR in Table E. For the MCQA probes, we report both their proportion of plurarity agreement and invariant accuracy. If the achieved proportion is far higher than the stated random probability on any probe, the model is considered to have multiple choice symbol binding proficiency. Table E shows the full results on CTR. Table E and Table E shows full results on the easy and hard variations of CL, respectively. Table E and Table E shows full results on the easy and hard variations of SI, respectively.

### v2

@@ Abstract
State-of-the-art large language models (LLMs) have demonstrated impressive code generation capabilities but struggle with real-world software engineering tasks, such as revising source code to address code reviews, hindering their practical use. Code review comments are often implicit, ambiguous, and colloquial, requiring models to grasp both code and human intent. This challenge calls for evaluating large language models’ ability to bridge both technical and conversational contexts. While existing work has employed the automated code refinement (ACR) task to resolve these comments, current evaluation methods fall short, relying on text matching metrics that provide limited insight into model failures and remain susceptible to training data contamination. To address these limitations, we introduce a novel evaluation benchmark, CodeReviewQA that enables us to conduct fine-grained assessment of model capabilities and mitigate data contamination risks. In CodeReviewQA, we decompose the generation task of code refinement into three essential reasoning steps: change type recognition (CTR), change localisation (CL), and solution identification (SI). Each step is reformulated as multiple-choice questions with varied difficulty levels, enabling precise assessment of model capabilities, while mitigating data contamination risks. Our comprehensive evaluation spans 72 recently released large language models on 900 manually curated, high-quality examples across nine programming languages. Our results show that CodeReviewQA is able to expose specific model weaknesses in code review comprehension, disentangled from their generative automated code refinement results.  [^1] Data Availability: https://huggingface.co/datasets/Tomo-Melb/CodeReviewQA
@@
Figure 1: CodeReviewQA results (%) of the top performing model (per scale class) based on ACR Accuracy. ACR: Automated Code Refinement, CTR: Change Type Recognition, CL: Change Localisation, SI: Solution Identification, E: Easy, H: Hard.
@@ 5.2 Model Selection Criteria
In total, we selected 72 state-of-the-art open source LLMs, that have satisfied the three criteria. The included models are considered state-of-the-art as of March, 2025. See Table 7 in Appendix F for descriptions of all 72 models. The models are grouped into five scales based on their model parameters: $\leq$3B, $\leq$9B, $\leq$16B, $\leq$34B, and $\leq$72B. We selected models with $\leq$72B parameters, as it is the largest non-quantised model class that we can support locally to extract answer probabilities.
@@ 6 Results Overview
To compare how models of different scales perform on CodeReviewQA, we conducted experiments using all 72 models. Due to space limitations, complete results on ACR and all three MCQA probes are presented in Appendix F. In Table 6, we compare the MCQA probe results of the top-2 models in terms of ACR from each scale class.
@@ 7 Insights from Probing ACR Failures
The results are presented in Table 7, including the percentage of failure on each of the MCQA probes,  [^3] For CL and SI, a failure in either difficulties is counted. and the overall failure on at least one probe ($\geq 1$ Probe). We find that all five models failed at least one MCQA probe for 76.5%-99.8% of the failure cases, indicating that the model struggled at the code review comprehension stage. CTR failures are seldom associated with the ACR failures, as this capability only accounts for 22.2%-37.4% of cases. In contrast, CL failures are often associated with ACR failures, particularly for the three smaller models, where 74.0%-99.4% of the cases overlap with ACR failures. SI failures account for 95.8% of cases for Qwen2.5-Coder-3B-Instruct. Thus, for the smallest model, most non-exact matches are associated with failures in both CL and SI. For both Qwen2.5-Coder-7B-Instruct and Qwen2.5-Coder-14B-Instruct, CL failure is the major factor associated with ACR failure. For the larger models, gemma-2-27b-it and Llama-3.1-70B-Instruct, the probe failures are more evenly distributed, varying based on the model’s specific incompetencies.
@@
Table 3: MCQA Failure (%) in Non-Exact Match cases of the top performing model (per scale class) in ACR.
@@ 8 Evaluating Data Contamination
$\text{PPL}(\textbf{X})=\exp(-\frac{1}{t}\sum^{t}_{t=0}\log p_{\theta}(x_{i}|x_{<i}))$ (5)
@@
$\text{NG}(\textbf{X})=\frac{1}{\eta\cdot K}\sum^{\eta}_{i=0}\sum^{K}_{j=0}I(X_{s_{j}:s_{j}+n},\hat{X}_{s_{j}:s_{j}+n})$ (6)
@@ Appendix D Descriptive Statistics
Specification ratio is the code edit distance of the change divided by the length of its respective code review comment. It is calculated as Code edit distanceComment $\frac{\text{Code edit distance}}{\text{Comment length}}$, the number of character edits with respect to each word in the comment. Since code review comments may be under-specified and implicit, we use specification ratio as a heuristic metric that incorporates this notion. Intuitively, examples with larger code edit distances and shorter comment lengths i.e. larger specification ratios, may be under-specified in its description of the required code change. Table D shows that the average specification ratio is 4.88. See Figure 11 for examples of different specification ratios.
Figure 3: Zero-Shot Prompt Templates.
Change Type Recognition (CTR) ### The following are multiple choice questions (with answers) that tests code review comprehension. Question: Given this Java code snippet, what type of change is the code review asking for? [Java] 1 private void launchZoomActivityAfterPermissionCheck(final View view) { 2 final Context ctx = view.getContext(); 3 final Intent zoomableIntent = new Intent(ctx, ZoomableActivity.class); 4 zoomableIntent.setData(Uri.parse(media.getImageUrl())); 5 zoomableIntent.putExtra("Origin", "MediaDetail"); 6 ctx.startActivity( 7 zoomableIntent 8 ); [/Java] [CODE REVIEW] Could you please rename to "MediaDetails"? [/CODE REVIEW] ### Possible answers: A. only add new lines of code B. only delete existing lines of code C. modify the code ### Answer with the letter symbol only. Answer: C Question: Given this CSharp code snippet, what type of change is the code review asking for? [CSharp] 1 public static class DateTimeDefinitions 2 public const string QuarterTypeRegex = @"(trimestral(mente)?)\$"; 3 public const string SemiAnnualTypeRegex = @"(semestral(mente)?)\$"; 4 public const string YearTypeRegex = @"(anual(mente)?)\$"; 5 public static readonly IList<string> FutureTerms = new List<string> 6 { 7 @"esea" 8 }; 9 } 10 } 11\ No newline at end of file [/CSharp] [CODE REVIEW] There are two issues here. One, this term is not a "future" term, so we need a better name. Second, this value is incorrect in PT, it should be "esse", "essa", "este", "esta". [/CODE REVIEW] ### Possible answers: A. only add new lines of code B. only delete existing lines of code C. modify the code ### Answer with the letter symbol only. Answer: C Question: Given this {lang} code snippet, what type of change is the code review asking for? [{lang}] {code_snippet} [/{lang}] [CODE REVIEW] {code_review} [/CODE REVIEW] ### Possible answers: A. only add new lines of code B. only delete existing lines of code C. modify the code ### Answer with the letter symbol only. Answer:
Figure 4: Few-Shot CTR Prompt Template.
Change Localisation (CL) {change_type} = add new lines of code under/delete code/modify code ### The following are multiple choice questions (with answers) that tests code review comprehension. Question: Given this Java code snippet, which line numbers is the code review asking to modify code? [Java] 1 private void launchZoomActivityAfterPermissionCheck(final View view) { 2 final Context ctx = view.getContext(); 3 final Intent zoomableIntent = new Intent(ctx, ZoomableActivity.class); 4 zoomableIntent.setData(Uri.parse(media.getImageUrl())); 5 zoomableIntent.putExtra("Origin", "MediaDetail"); 6 ctx.startActivity( 7 zoomableIntent 8 ); [/Java] [CODE REVIEW] Could you please rename to "MediaDetails"? [/CODE REVIEW] ### Possible answers: A. 3 B. 4 C. 5 D. 6 ### Answer with the letter symbol only. Answer: C Question: Given this CSharp code snippet, which line numbers is the code review asking to modify code? [CSharp] 1 public static class DateTimeDefinitions 2 public const string QuarterTypeRegex = @"(trimestral(mente)?)\$"; 3 public const string SemiAnnualTypeRegex = @"(semestral(mente)?)\$"; 4 public const string YearTypeRegex = @"(anual(mente)?)\$"; 5 public static readonly IList<string> FutureTerms = new List<string> 6 { 7 @"esea" 8 }; 9 } 10 } 11\ No newline at end of file [/CSharp] [CODE REVIEW] There are two issues here. One, this term is not a "future" term, so we need a better name. Second, this value is incorrect in PT, it should be "esse", "essa", "este", "esta". [/CODE REVIEW] ### Possible answers: A. 5, 7 B. 2, 3 C. 1, 4 D. 8, 11 ### Answer with the letter symbol only. Answer: A Question: Given this {lang} code snippet, which line numbers is the code review asking to {change_type}? [{lang}] {code_snippet} [/{lang}] [CODE REVIEW] {code_review} [/CODE REVIEW] ### Possible answers: A. {option_a} B. {option_b} C. {option_c} D. {option_d} ### Answer with the letter symbol only. Answer:
Figure 5: Few-Shot CL Prompt Template.
Solution Identification (SI) ### The following are multiple choice questions (with answers) that tests code review comprehension. Question: Given this Java code snippet, which code revision is the code review asking for? [Java] 1 private void launchZoomActivityAfterPermissionCheck(final View view) { 2 final Context ctx = view.getContext(); 3 final Intent zoomableIntent = new Intent(ctx, ZoomableActivity.class); 4 zoomableIntent.setData(Uri.parse(media.getImageUrl())); 5 zoomableIntent.putExtra("Origin", "MediaDetail"); 6 ctx.startActivity( 7 zoomableIntent 8 ); [/Java] [CODE REVIEW] Could you please rename to "MediaDetails"? [/CODE REVIEW] ### Possible answers: A. 5- zoomableIntent.putExtra("Origin", "MediaDetail"); { 5+ zoomableIntent.postExtra("Origin", "MediaDetails"); B. 5- zoomableIntent.putExtra("Origin", "MediaDetail"); { 5+ zoomableIntent.getExtra("Origin", "MediaDetails"); C. 5- zoomableIntent.putExtra("Origin", "MediaDetail"); { 5+ zoomableIntent.putExtra("Origin", "MediaDetails"); D. 5- zoomableIntent.putExtra("Origin", "MediaDetail"); { 5+ zoomableIntent.renameExtra("Origin", "MediaDetails"); ### Answer with the letter symbol only. Answer: C Question: Given this CSharp code snippet, which code revision is the code review asking for? [CSharp] 1 public static class DateTimeDefinitions 2 public const string QuarterTypeRegex = @"(trimestral(mente)?)\$"; 3 public const string SemiAnnualTypeRegex = @"(semestral(mente)?)\$"; 4 public const string YearTypeRegex = @"(anual(mente)?)\$"; 5 public static readonly IList<string> FutureTerms = new List<string> 6 { 7 @"esea" 8 }; 9 } 10 } 11\ No newline at end of file [/CSharp] [CODE REVIEW] There are two issues here. One, this term is not a "future" term, so we need a better name. Second, this value is incorrect in PT, it should be "esse", "essa", "este", "esta". [/CODE REVIEW] ### Possible answers: A. 5- public static readonly IList<string> FutureTerms = new List<string> 5+ public static readonly IList<string> ThisTerms = new List<string> 7- @"esea" 7+ @"esse" 8+ @"essa" 9+ @"este" 10+ @"esta" B. 5- public static readonly IList<string> FutureTerms = new List<string> 5+ public static readonly IList<string> NotTerms = new List<string> 7- @"esea" 7+ @"esse" 8+ @"essa" 9+ @"este" 10+ @"esta" C. 5- public static readonly IList<string> FutureTerms = new List<string> 5+ public static readonly IList<string> futureTerms = new List<string> 7- @"esea" 7+ @"esse" 8+ @"essa" 9+ @"este" 10+ @"esta" D. 5- public static readonly IList<string> FutureTerms = new List<string> 5+ public static readonly IList<string> betterTerms = new List<string> 7- @"esea" 7+ @"esse" 8+ @"essa" 9+ @"este" 10+ @"esta" ### Answer with the letter symbol only. Answer: A Question: Given this {lang} code snippet, which code revision is the code review asking for? [{lang}] {code_snippet} [/{lang}] [CODE REVIEW] {code_review} [/CODE REVIEW] ### Possible answers: A. {option_a} B. {option_b} C. {option_c} D. {option_d} ### Answer with the letter symbol only. Answer:
Figure 6: Few-Shot SI Prompt Template.
[…]
Figure 7: Examples of Variation in Difficulty.
Figure 8: Examples of Different Comment Lengths.
Figure 9: Examples of Different Code Edit Distances.
Figure 10: Examples of Different Code Element Ratios.
Figure 11: Examples of Different Specification Ratios.
Appendix E Implementation Details
The experiments were carried out on a single node consisting of 64 cores (Intel(R) Xeon(R) Platinum 8462Y+ @ 2.80GHz), 928GBs of RAM and four GPUs (NVIDIA H100-80GB SXM5). For efficient inference, we utilised the default vLLM implementation with Hugging Face models. 444 https://huggingface.co/docs/hub/en/models-the-hub For all tasks, we set the temperature to zero for greedy decoding.
Appendix F Experimental Results on 72 models
This section shows our complete results. The details of all evaluated LLMs are presented in Table 7, and their achieved exact match rate on ACR in Table G. For the MCQA probes, we report both their proportion of plurarity agreement and invariant accuracy. If the achieved proportion is far higher than the stated random probability on any probe, the model is considered to have multiple choice symbol binding proficiency. Table G shows the full results on CTR. Table G and Table G shows full results on the easy and hard variations of CL, respectively. Table G and Table G shows full results on the easy and hard variations of SI, respectively.
Appendix G Advanced Prompting
Since our main results were based on zero-shot prompting, we further investigated the effects of advanced prompting i.e., few-shot, chain-of-thought, on MCQA probe performance. Specifically, we were interested if advanced prompting could increase the performance of the top performing model i.e., Llama-3.1-70B-Instruct. For few-shot prompting, we considered both one-shot and two-shot scenarios with real examples held out from the benchmark. For the few-shot prompt templates, see Figure 4 for CTR, Figure 5 for CL and Figure 6 for SI. For chain-of-thought, we only considered zero-shot due to the lack of real reasoning traces to use as examples. This was invoked by appending "Let’s think step by step" to the end of the zero-shot prompts and instructing the model to finalise their answer with "The final answer is".
[…]
Zero-Shot 68.4 74.7 69.0 84.2 76.7
One-Shot 74.1 68.6 61.9 59.9 55.6
Two-Shot 76.9 75.8 68.7 57.3 54.6
Chain-of-Thought 68.6 65.8 60.1 66.6 60.0
[…]
CL: Change Localisation, SI: Solution Identification
Table 6: MCQA Probe Accuracy (%) of Llama-3.1-70B-Instruct w/ Advanced Prompting.
The results are shown in Table G. Overall, we find that advanced prompting does not outperform the zero-shot strategy. For CTR, we find that zero-shot performed similar to chain-of-thought, where both accuracies were approximately 68%. In contrast, the few-shot methods could achieve 74.1-76.9%, where two-shot prompting was the top performing prompt strategy. In terms of CL easy, we find that one-shot and chain-of-thought were the worst performers, achieving 68.6% and 65.8%, respectively. In contrast, zero-shot and two-shot performed similarly, achieving 74.7% and 75.8%, respectively. For the remaining MCQA probes, all advanced prompting techniques underperformed against the zero-shot strategy, especially for SI.
Whilst chain-of-thought consistently degraded the model’s performance on the MCQA probes, few-shot prompting showed promise. Specifically, two-shot prompting was the top performing strategy in two of the five probes, without optimising for the best examples. We did not optimise this selection due to the lack of clean code review examples. As a result, we only used two held out examples collected during the manual curation process. Future researchers can investigate the effects of advanced example selection e.g., retrieval-augmented generation, for few-shot prompting strategies.